OpenStudy (anonymous):

what is the phase shift of the function y=1/2sin(x/3-pi)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

y = A*sin(Bx - C) + D has a phase shift of w = C/B

OpenStudy (anonymous):

1/2?

jimthompson5910 (jim_thompson5910):

B = ??? C = ??

OpenStudy (anonymous):

B=x/3 C=-pi

jimthompson5910 (jim_thompson5910):

no B = 1/3 and C = pi

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

so what do you get

OpenStudy (anonymous):

pi?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

w = C/B w = pi/(1/3) w = ???

OpenStudy (anonymous):

pi/3

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (anonymous):

-pi?

OpenStudy (anonymous):

i multiply by 3/1 right?

jimthompson5910 (jim_thompson5910):

w = pi/(1/3) w = (pi/1)/(1/3) w = (pi/1)*(3/1) w = ???

OpenStudy (anonymous):

3pi?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

you see i had that originally, but I'm confused about it because its not one of the choices

jimthompson5910 (jim_thompson5910):

what are your choices

OpenStudy (anonymous):

-pi, 1/2, pi, pi/3

jimthompson5910 (jim_thompson5910):

are you sure that they don't want the amplitude?

OpenStudy (anonymous):

positve, they don't want the amplitude they want the phase shift

jimthompson5910 (jim_thompson5910):

hmm well the closest thing I can think of is pi since C = pi, but the phase shift usually involves B

OpenStudy (anonymous):

I guess ill just ask my teacher about it tomorrow, thank you for your effort though, much appreciated!

jimthompson5910 (jim_thompson5910):

you're welcome